Find the measures of variability of the given table. A nutritionist thinks that people with low income average have less Recommended Daily Allowance (RDA) of 800mg intakes of Calcium (standard unit of Calcium for adult). He got a sample of 40 people with income below poverty level. The results are obtained in the following frequency distribution. Intake (mg) Frequency 101-200 1 201-300 1 301-400 9 401-500 13 501-600 10 601-700 6
1
Expert's answer
2020-10-13T19:53:11-0400

Solution


The measures of variability are:

Range, Interquartile range, variance, standard deviation and mean deviation.




Range:

Largest value - Smallest value


"=700-101=599"

Range is 599


Interquartile Range

"={Q_3-Q_1}"

"Q_1= L+{h \over f}({n\over4}-cf)""Q_3= L+{h \over f}({3n\over4}-cf)"

where L= lower limit of the class, h= class width, f= frequency of the class, n= no. of samples, cf= cumulative frequency above the quartile class

"{1 \over 4}*40=10 \therefore" Q1 lies on the third class between 301-400.


"\therefore Q_1=300.5+{100 \over 9}({40 \over 4} -2) = 389.38"

"{3 \over 4}*40=30 \therefore Q_3" lies on the fifth class between 501-600

"\therefore Q_3=500.5+{100 \over 10}({3*40 \over 4} -24) = 560.5"

"IQR={560.5-389.38}=171.12"

Inter-quartile range is 85.56


Variance

"var={\sum x^2f -({\sum fx \over n})^2 \over n-1}"

"= ({9398810-{18820 \over 40} \over 39})=240983.0641"

Variance is 240983.0641


Standard deviation

"SD = \sqrt{var} = \sqrt{240983.0641}""=490.9002588"

Standard deviation is 490.9003


Mean deviation

"={\sum f(x- \bar{x}) \over n} = {0 \over 40}=0"